# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

### v1

@@ Abstract
The success of DNNs often depends on training with large-scale datasets, but building such datasets is both expensive and challenging. Consequently, public datasets from open-source platforms like HuggingFace have become popular, posing significant risks of data poisoning attacks. Existing backdoor defenses in NLP primarily focus on identifying and removing poisoned samples; however, purifying a backdoored model with these sample-cleaning approaches typically requires expensive retraining. Therefore, we propose Greedy Module-Substitution (GMS), which identifies and substitutes “deadwood” modules (i.e., components critical to backdoor pathways) in a backdoored model to purify it. Our method relaxes the common dependency of prior model purification methods on clean datasets or clean auxiliary models. When applied to RoBERTa-large under backdoor attacks, GMS demonstrates strong effectiveness across various settings, particularly against widely recognized challenging attacks like LWS, achieving a post-purification attack success rate (ASR) of 9.7% on SST-2 compared to 58.8% for the best baseline approach.
Cut the Deadwood Out: Post-Training Model Purification
with Selective Module Substitution
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various domains Krizhevsky et al. (2012); Bowman et al. (2016), largely driven by training on large-scale datasets. Given the high cost of collecting and labeling data from scratch, incorporating public datasets from open-source platforms like HuggingFace in training has become a common practice, particularly in NLP Howard and Ruder (2018); Ding et al. (2023). However, this growing reliance on open-source datasets from unaudited platforms raises significant security concerns Yang et al. (2021a); Zhang et al. (2021), especially for commercial use. Malicious actors can exploit this ecosystem by publishing poisoned datasets Xu et al. (2021) to embed backdoors into models for potential benefit. These backdoors allow attackers to manipulate models, causing targeted misbehavior when triggered by specific inputs while operating normally on clean inputs. As a result, they often remain stealthy until their triggers are activated.
Figure 1: Parameter-level editing (left) prunes parameters "most relevant" to backdoor features but leaves many backdoor-related parameters intact within each module, allowing backdoor pathways to potentially persist. In contrast, our module-level substitution (right) effectively disrupts these pathways by completely replacing backdoor-connected modules with counterparts from a proxy model (not required to be clean).
Existing defense strategies against backdoor attacks in NLP primarily focus on backdoor samples detection, such as detecting and removing poisoned samples leveraging their distinct characteristics compared to clean ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2022, 2023); Xu et al. (2024). However, their effectiveness is often suboptimal Min et al. (2024), especially as backdoor triggers evolve to be stealthier and harder to detect or reverse Qi et al. (2021b). Moreover, while these methods follow a fine-tuning or adversarial training for unlearning Wei et al. (2023) to sanitize poisoned models, they cannot effectively purify a backdoored model without incurring a significant cost of retraining Wu et al. (2022). Then a practical question arises: How can we cost-efficiently eliminate the backdoor features from an existing model after users reported some compromise?
While only a few model purification defenses in NLP exist, they predominantly focus on parameter-level editing. For instances, Fine-mixing Zhang et al. (2022) and Fine-purification Zhang et al. (2023) mix a clean pre-trained language model (PLM) with a backdoored model and rely on fine-tuning with a clean dataset for purification. Similarly, Zhao et al. (2024) prunes attention heads potentially affected by poisoned data and fine-tunes the pruned model using a clean dataset. However, recent studies have observed that backdoor features are encoded as complex pathways in victim models, with significant overlap between backdoor and clean neurons at the module level Jebreel et al. (2023); Li et al. (2024). As illustrated in Figure 1 (Left), parameter-level editing will leave hidden vulnerabilities in the non-pruned connections, thus making the defense effectiveness heavily relies on fine-tuning with a clean dataset to restore clean features Liu et al. (2018), an assumption that is both strong and rarely feasible in practice.
Here, we develop a method that effectively disrupts the model’s backdoor pathways without relying on the assumption of access to a clean dataset. Note that the abundance of public platforms for open-source models (e.g., HuggingFace) offers semantically similar yet independent models for any task domain Semwal et al. (2018). Therefore, we propose a simple yet effective module-level purification defense, Greedy Module-Substitution (GMS), which identifies critical modules for backdoor propagation and purifies them by replacing the parameters with those of a proxy model. While this approach eliminates the need for a clean dataset to restore semantic knowledge, the proxy model is not required to be clean either, since the pathways of two different backdoored models are less likely to be connected (empirically verified in Table 5). The primary contributions of this work are as follows.
@@
    We propose the first efficient (quadratic time complexity) module-level backdoor purification defense (GMS), which relaxes the strong reliance on clean auxiliary models and clean datasets required by prior defenses, as shown in Table 5 and Table 6.
@@
    Our method significantly outperforms baseline approaches against challenging backdoor attacks like LWS and HiddenKiller, e.g., on SST-2 victim model, GMS reduces the ASR on LWS to 9.7%, compared to 12.8% for the benign model and 58.8% for the best baseline.
@@
    Our GMS purification strategy exhibits strong transferability across different attacks. Even the strategy with few module substitutions can effectively defend against other backdoor attacks, making it suitable when the victim dataset is inaccessible (Section 5.3 and Table 7).
@@ 2 Related works
Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present. They can be categorized into two series Wu et al. (2022): (1) Data-poisoning attacks, exemplified by BadNets Gu et al. (2017), involve tampering with a subset of training data by adding triggers and altering labels, making them practical in real-world scenarios with minimal attacker assumptions Turner et al. (2019); Carlini and Terzis (2022); Goldblum et al. (2023); Li et al. (2021c). Initially explored in image classification, backdoor attacks have since raised significant security concerns in NLP Dai et al. (2019); Qi et al. (2021d, c) with triggers range from misspelled words Chen et al. (2021), rare words Yang et al. (2021a); Kurita et al. (2020) and syntactic structures Qi et al. (2021c) to text styles Pan et al. (2022), posing serious challenges for detection. (2) Training-control attacks, on the other hand, assume complete control over the training process and data Nguyen and Tran (2021); Kurita et al. (2020), with advances such as layer-wise weight poisoning ensuring backdoor persistence even after fine-tuning Li et al. (2021a). In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
Backdoor defense
@@
Backdoor defense approaches can be categorized according to their strategies: (1) Backdoor sample detection focuses on identifying poisoned samples in the dataset. The defense approaches usually include detecting and removing backdoor samples, leveraging the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2019, 2022, 2023); Xu et al. (2024). These strategies are typically applied during the training phase or involve fine-tuning after removing backdoor samples. (2) Backdoor model purification: These methods aim to eliminate backdoor features from a well-trained model. Common approaches involve merging parameters with other proxy models Arora et al. (2024); Chen et al. (2024a), or fine-tuning or pruning using a limited amount of clean data Wu and Wang (2021); Min et al. (2023); Liu et al. (2018); Zhao et al. (2024); Zhu et al. (2023). These techniques not only reduce training costs but also achieve state-of-the-art defense performance Zhu et al. (2023); Zhao et al. (2024); Arora et al. (2024). Our work advances the second line of research by minimizing the number of auxiliary models to one and reducing dependence on the clean dataset.
@@ 3 Preliminary
    Layer Set ($\mathcal{L}$) is the set of block indices, capturing the architecture’s depth and structure.
@@
For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules. Attention modules contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively. The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$. For simplicity and consistency, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper. Accordingly, we define the module set $\mathcal{M}$ for Transformer models as $\{Q,K,V,O,F,P\}$ and layer set =$\mathcal{L}=\{1,2,\ldots,L\}$.
Evaluation metrics
Following previous literature (Qi et al., 2020), we adopt two metrics: (1) Clean Accuracy (CACC) measures the prediction accuracy on clean samples, and (2) Attack Success Rate (ASR) measures the attack accuracy on poisoned test set. A higher CACC reflects better utility, while lower ASR indicates stronger defense.
Figure 2: The pipeline of our method. Step 1: Extract two small proxy datasets from the suspected dataset $\mathcal{D}$ used for computing score in Equation 1. Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation. Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$. For more details, refer to Algorithm 1.
@@ 4 Methodology
In Section 4.1, we introduce the problem setting, and in Section 4.2 we present our methods.
@@ 4.1 Problem setting
Threat Model
We consider the scenario of data poisoning-based backdoor attacks, where an attacker releases a poisoned dataset $\mathcal{D}$ to embed backdoors into any model trained on this dataset. Starting with a clean dataset $\mathcal{D}_{\text{cl}}=\{(x_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}\subseteq\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ denote the input and label spaces, respectively, the attacker selects a subset $\mathcal{D}_{\text{sub}}\subset\mathcal{D}_{\text{cl}}$, and then modified each sample $(x,y)\in\mathcal{D}_{\text{sub}}$ into a poisoned sample $(\tilde{x},y_{t})$ where $\tilde{x}=g(x,\Delta)$, the pre-defined trigger is $\Delta$, $g$ is a generation function, and $y_{t}$ is the target label. The poisoned subset is $\mathcal{D}_{\text{poi}}=\{(\tilde{x},y_{t})\mid(x,y)\in\mathcal{D}_{\text{sub}}\}$. Finally, the attackers publish the poisoned dataset constructed as $\mathcal{D}=(\mathcal{D}_{\text{cl}}\setminus\mathcal{D}_{\text{sub}})\cup\mathcal{D}_{\text{poi}}$.
Defense Scenario
The defender (e.g., a model vendor) has already trained a (potentially backdoored) victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$. While the defender has access to $\mathcal{D}$, they are unaware of all possible backdoor patterns. [^1] As shown in Section 4.2.2, the full training dataset is not strictly required; a small subset of poisoned data reported by users or automatically detected by algorithms is sufficient for applying our defense. The defender has white-box access to a trained victim model $M_{v}$ and a proxy model $M_{p}$ trained on similar downstream tasks.
@@ 4.2 Our method
We aim to disrupt the backdoor connection in victim model $M_{v}$ while preserving the knowledge required for the original task. Models trained on similar tasks tend to capture similar domain-specific knowledge Semwal et al. (2018). Hence, we replace backdoor-connected modules in the $M_{v}$ with counterparts from a proxy model $M_{p}$ that shares similar domain knowledge. Notably, $M_{p}$ does NOT need to be benign, as even a malicious proxy model lacking identical backdoor pathways will fail to restore the disrupted backdoor connection.
Our main algorithm, Greedy Module-Substitution (GMS), is detailed in Section 4.2.3. Section 4.2.1 defines the objective optimized by GMS, and Section 4.2.2 describes the method to construct the proxy datasets used in evaluation.
@@ 4.2.1 Objective
The defender’s objective is to identify a purified model $M_{pure}$ from all possible substituted models $M$ that effectively eliminates the backdoor while preserving task utility. This can be formulated as
@@
where $\delta_{\text{asr}}(M)$ measures the extent of backdoor removal by quantifying the reduction in ASR and $\delta_{\text{acc}}(M)$ assesses the impact on task utility by measuring the decrease in CACC. However, since the defender lacks access to the clean dataset and poisoned test set, we adopt the method described in Section 4.2.2 to extract two proxy datasets, $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, from the suspected dataset $\mathcal{D}$ for ASR and CACC estimation.
The goal is to maximize $\delta_{\text{asr}}(M)$ while minimizing $\delta_{\text{acc}}(M)$. The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor defense and task utility. A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
In this section, we introduce the process of extracting two proxy datasets from the suspected dataset $\mathcal{D}$: a proxy clean set $\mathcal{D}_{clean}$, which is identified with high probability as clean (i.e., they are likely to be clean though not guaranteed to be free from poisoning), and a proxy poison set $\mathcal{D}_{poison}$, which is identified with high confidence as being poisoned. Both proxy datasets are not required to be purely clean or poisoned. When the defender has no prior knowledge of the poisoned samples, various methods can be used to construct proxy datasets, such as sample diagnosis Gao et al. (2019); Huang et al. (2022); Guo et al. (2023) and trigger inversion Wang et al. (2019, 2022, 2023); Xu et al. (2024). If additional prior knowledge, such as a user-reported clean dataset, is available (as tested in Table 6), the performance of our method is expected to be improved. For simplicity, in this paper, we use random sampling to construct the proxy clean set $\mathcal{D}_{clean}$, based on the practical assumption that the poison samples will not dominate a training set. To construct the proxy poison set $\mathcal{D}_{poison}$, we adopt an existing backdoor sample detection method from SEEP He et al. (2024), which leverages an observation that backdoored models tend to exhibit over-confidence on poisoned samples Li et al. (2021b). SEEP utilizes training dynamics to identify anomalous samples in the training corpus, which are proven to have a high probability of being backdoor samples. In our example, we only focus on the most confident samples extracted (i.e., seed set). In Section 5.2, we show that extracting only 1% of the original poisoned data in $\mathcal{D}_{poison}$ and allowing even 20% poisoned data in $\mathcal{D}_{clean}$ are sufficient to launch an effective defense using our Greedy-Module-Substitution (GMS).
4.2.3 Greedy search for purified model $M_{pure}$
Using proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$, we can estimate ASR and ACC to quantify the level of backdoor defense and task utility for a model $M$, we define the following two metrics:
@@ 4.2.3 Greedy search for purified model $M_{pure}$
Here, $s_{acc}(M)$ represents the CACC of $M$ on $\mathcal{D}_{clean}$, while $s_{asr}(M)$ denotes its ASR on $\mathcal{D}_{poison}$. Similarly, $s_{acc}(M_{v})$ and $s_{asr}(M_{v})$ are the ASR and CACC for the original victim model $M_{v}$. In this context, $\delta_{\text{asr}}(M)$ quantifies the reduction in ASR achieved by a purified model $M$, while $\delta_{\text{acc}}(M)$ reflects the drop in task utility. Plugging them back into the objective (Equation 1) gives us an score. [^2] An inaccurate $\mathcal{D}_{poison}$ may lead to a smaller $\delta_{\text{asr}}$, while an impure $\mathcal{D}_{clean}$ might result in a larger $\delta_{\text{acc}}(M)$. However, since the metrics are always calculated on fixed proxy datasets, these effects can be controlled by adjusting $\alpha$.
While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\cdot 12}$ for a BERT-base (Devlin et al., 2019)). Thus, we propose a greedy algorithm that transforms the problem into a feasible search by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
Algorithm 1 Greedy-Module-Substitution (GMS)

### v2

# Cut the Deadwood Out: Backdoor Purification via Guided Module Substitution

@@ Abstract
Model NLP models are commonly trained (or fine-tuned) on datasets from untrusted platforms like HuggingFace, posing significant risks of data poisoning attacks. A practical yet underexplored challenge arises when such backdoors are discovered after model deployment, making retraining-required defenses less desirable due to computational costs and data constraints. In this work, we propose Guided Module Substitution (GMS), an effective retraining-free method based on guided merging of the victim model with just a single proxy model. Unlike prior ad-hoc merging defenses, GMS uses a guided trade-off signal between utility and backdoor to selectively replaces modules in the victim model. GMS offers four desirable properties: (1) robustness to the choice and trustworthiness of the proxy model, (2) applicability under inaccurate data knowledge, (3) stability across hyperparameters, and (4) transferability across different attacks. Extensive experiments on encoder models and decoder LLMs demonstrate the strong effectiveness of GMS. GMS significantly outperforms even the strongest defense baseline, particularly against challenging attacks like LWS. The code is available at https://github.com/weijun-l/guided-module-substitution.
Cut the Deadwood Out: Backdoor Purification
via Guided Module Substitution
@@ 1 Introduction
Modern NLP models are frequently adapted to specific downstream tasks through fine-tuning on custom datasets Howard and Ruder (2018). In practice, these datasets are often collected from diverse sources, some of which may be unreliable (e.g., open repositories like HuggingFace Lhoest et al. (2021)). As a result, models can unknowingly incorporate poisoned data, leading to backdoor vulnerability Zhang et al. (2021); Xu et al. (2021): a backdoored model, trained with poisoned data, behaves normally on clean inputs but misbehaves when exposed to the trigger Bai et al. (2025).
A critical yet underexplored problem arises when such backdoors are discovered after model deployment. Existing defenses, such as data filtering and retraining (He et al., 2024), pruning followed by fine-tuning Liu et al. (2018); Zhao et al. (2024b), or unlearning backdoors based on clean Li et al. (2023) or poisoned data Min et al. (2024); Li et al. (2021b); Chen et al. (2022), typically incur substantial computational costs Wu et al. (2022)–especially for large-scale models or proprietary pipelines. This raises a practical and timely question: Can we effectively purify a trained backdoored model without retraining?
Recent work (Arora et al., 2024) has proposed model merging as a cost-efficient and retraining-free defense method for model purification: leveraging the abundance of open-source proxy models online [^1] For reference, on Hugging Face, there are 74 LLaMA-7B and 262 BERT-base models trained on SST-2 alone, and the pool of proxy models is even larger when considering all datasets within the same task domain., this approach merges several proxy models–backdoored or not–with the victim model. While promising, this approach relies on ad hoc merging of multiple models with no principled selection, such as utility and defense performance. Its effectiveness is thus unpredictable and sensitive to the number of models involved. Our later experiments in Section˜5.2 demonstrate that when only a single proxy model is available, such defenses become ineffective. More seriously, In practice, collecting more homogeneous proxy models also requires more effort and often harm utility more Zhou et al. (2025a); Wang et al. (2025b), even though more merged parameters can help neutralize backdoors. We therefore target a more challenging scenario–designing guided merging using only one proxy model.
In this work, we propose a simple yet effective alternative: guided merging with a single proxy model. Rather than blindly merging models, we explicitly guide the merging process using a signal that approximates the trade-off between utility and backdoor risk. Intuitively, if retaining a particular module from the victim model contributes significantly more to the attack performance than to task utility, it is likely critical for encoding backdoor features. We thus progressively identify such modules and replace them with their counterparts from the proxy model, as long as the resulting utility degradation remains limited relative to backdoor mitigation gain. We refer to this method as Guided Module-Substitution (GMS). While training-free and more efficient than prior model-merging-based defenses, our method GMS not only retains the robustness of model-merging defenses to proxy model trustworthiness, but is also tolerant of imperfect data knowledge, stable across hyperparameters, and exhibits strong transferability across different attacks. Further discussion of these desirable properties is in Section˜5.3.
The contributions of this work are as follows.
@@
    We propose GMS, the first backdoor purification method based on guided merging of a victim model with a single proxy model. This design enables purification without retraining while remaining efficient and practical.
@@
    Experiments on standard NLP models and LLMs (Section˜5.2) show that GMS consistently surpasses competitive defense baselines, with especially strong performance against the more challenging attacks like LWS and HiddenKiller.
@@
    We further validate four desirable properties of GMS in Section˜5.3 and Sections˜B.4, B.5, B.6, B.7, B.8, B.9, B.10, B.11 and B.12, namely (1) robustness to proxy model choices, (2) applicability to varied data conditions, (3) stability across hyperparameters, and (4) transferability across attacks.
@@ 2 Related works
Backdoor attacks aim to manipulate models to behave normally on benign inputs while exhibiting attacker-controlled behavior when specific triggers are present. They can be categorized into two series Wu et al. (2022): (1) Data-poisoning attacks, exemplified by BadNets Gu et al. (2017), involve tampering with a subset of training data by adding triggers and altering labels, making them practical in real-world scenarios with minimal attacker assumptions Turner et al. (2019); Carlini and Terzis (2022); Goldblum et al. (2023); Li et al. (2021c). Initially explored in image classification, backdoor attacks have since raised significant security concerns in NLP Dai et al. (2019); Qi et al. (2021c, b) with triggers range from misspelled words Chen et al. (2021), rare words Yang et al. (2021a); Kurita et al. (2020) and syntactic structures Qi et al. (2021b) to text styles Pan et al. (2022), posing serious challenges for detection. (2) Training-control attacks, on the other hand, assume complete control over the training process and data Nguyen and Tran (2021); Kurita et al. (2020), with advances such as layer-wise weight poisoning ensuring backdoor persistence even after fine-tuning Li et al. (2021a). In this work, we focus on data-poisoning backdoor attacks in NLP due to their widespread applicability and practical implications.
Backdoor defense.
@@
Post-training backdoor defense approaches can be broadly categorized as: (1) Backdoor sample detection followed by training, which first filters poisoned samples from the dataset and then retrains or fine-tunes the model. These defenses typically rely on the assumption that poisoned samples exhibit distinct characteristics compared to clean (benign) ones, enabling their detection Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), or attempt to reverse-engineer backdoor triggers to neutralize their impact Wang et al. (2019, 2022, 2023); Xu et al. (2024). Afterward, such strategies usually involve unlearning the identified backdoor samples or retraining/fine-tuning on the filtered dataset. However, their effectiveness is often unsatisfactory Sun et al. (2025); Wang et al. (2023); Qi et al. (2021a), primarily because accurately identifying all poisoned samples is (increasing) difficult Wu et al. (2022); Qi et al. (2021a); Yang et al. (2021b); Tran et al. (2018); He et al. (2023), leaving residual backdoor behavior even after retraining or unlearning. Our method, in this regard, compensates for these shortcomings by being more robust to data filtering quality, and thus provides a more reliable alternative to unlearning- or retraining-based defenses. (2) Backdoor model purification: These methods aim to eliminate backdoor features directly from a well-trained model and are generally regarded as achieving state-of-the-art defense performance Zhu et al. (2023); Zhao et al. (2024b). Common approaches involve merging parameters with other proxy models Arora et al. (2024); Chen et al. (2024a), or pruning and fine-tuning using a limited amount of clean data Wu and Wang (2021); Min et al. (2023); Liu et al. (2018); Zhao et al. (2024b); Zhu et al. (2023), which may sometimes be assisted by clean proxy models as well Liu et al. (2018); Zhang et al. (2023). Our work advances the second line of research by minimizing the number of auxiliary models to one and by eliminating the need for additional training and strictly clean datasets.
Model merge.
Model merging methods have emerged as an efficient way to build powerful models by combining existing trained domain-specialized models without requiring extensive retraining Wortsman et al. (2022); Ilharco et al. (2023); Cheng et al. (2025). Most advances focus on developing utility-preserving multi-task models (e.g., merging models trained on different tasks like coding and mathematics He et al. (2025)), or stronger domain-specific models by combining within-domain models Wortsman et al. (2022). Surprisingly, recent work shows that naive weight averaging across multiple models can mitigate backdoors Arora et al. (2024). Yet such blind merging is inefficient–requiring unpredictable trial-and-error in model collection, since one cannot know in advance how many models are needed for purification–and often harmful, as merging many models can yield utility drops from parameter interference and redundancy Zhou et al. (2025a); Wang et al. (2025b). Thus, our work address these issues by proposing effective and efficient guided merging with a single proxy model. Note that while there are some works studying how model merging systems themselves can be exploited for backdooring, leveraging the fact that merging methods aim to preserve (task) parameters from all models Wang et al. (2025a); Zhang et al. (2024); Yuan et al. (2025), our objective is fundamentally different: we use merging as a tool to purify a single target model, where only the target model’s clean task is relevant and other tasks can all be discarded. Since our main goal is to demonstrate guided merging can act as a defense–and obtaining a single homogeneous proxy is practical–we focus on the standard homogeneous setting. As heterogeneous merging itself remains an open challenge Xu et al. (2025); Du et al. (2025), exploring its use for defense is left to future work.
Safety Localization.
Recent studies suggest that safety-critical behaviors in LLMs are often localized to specific layers or components. For example, jailbreak defenses identify safety layers that are critical for aligning harmful queries Zhao et al. (2024a); Ouyang et al. (2025); Zhou et al. (2024); a small set of safety layers Li et al. (2025) or safety modules (e.g., attention heads) Zhou et al. (2025b) are crucial for distinguishing malicious from benign inputs; and modifying specific safety-related parameters can further enhance alignment Wei et al. (2024). Similarly, backdoor research shows that certain layers Jebreel et al. (2023) and even neurons or heads Zhao et al. (2024b) disproportionately influence attack success, enabling defenses through targeted editing or masking. These findings motivate our focus on module- and layer-level guided merging, which balances efficiency with effectiveness.
@@ 3 Preliminary
    Layer Set ($\mathcal{L}$) is the set of block indices that captures the depth of the model.
@@
For each Transformer block, the input will be processed through Attention and Feed-Forward Network (FFN) modules. An attention module contains four components: $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{O}$, projecting the input into query, key, value, and the final output representation, respectively. The following FFN contains two components: $W_{F}$ for the forward layer and $W_{P}$ for the projection layer, connected by a non-linear activation function.
In total, each Transformer block contains six key functional modules with parameters $W_{Q},W_{K},W_{V},W_{O},W_{F},$ and $W_{P}$. For simplicity, we adopt notations $Q,K,V,O,F,$ and $P$ to refer to these weight matrices throughout the paper. Accordingly, we define the module set for Transformer models $\mathcal{M}=\{Q,K,V,O,F,P\}$ and layer set $\mathcal{L}=\{1,2,\ldots,L\}$.
Figure 1: The pipeline of our method. Step 1: Extract two small proxy datasets used for computing score in Equation˜1. Step 2: Iteratively update a substitution matrix $S$ to greedily maximize the trade-off score between backdoor removal and utility preservation. Step 3: Return purified model $M_{pure}$ corresponding to the best substitution matrix $S_{best}$. For more details, refer to Algorithm˜1.
@@ 4 Methodology
In Section˜4.1, we introduce the problem setting, and we present our methods in Section˜4.2.
@@ 4.1 Problem setting
Defense Setting.
The defender (e.g., a model vendor) has trained a potentially backdoored victim model $M_{v}$ on an unauthenticated dataset $\mathcal{D}$, which may contain poisoned samples. Following the standard setting Li et al. (2021b), we assume the defender has full control over the training process but lacks prior knowledge of the backdoor patterns, their proportion, or their distribution within $\mathcal{D}$. If an isolation method is employed, it may only identify a subset of the poisoned examples, as perfect detection is not guaranteed. The defender has white-box access to the trained victim model $M_{v}$, as well as access to a proxy model $M_{p}$ trained on similar downstream tasks.
@@ 4.2 Our method
We aim to combine a victim model with a proxy model. Unlike traditional model merging Arora et al. (2024), which prioritizes utility across all models’ tasks, our method targets backdoor removal and requires preserving only the clean task performance of the target model. This enables a more targeted substitution strategy: directly replacing backdoor-encoded modules in the victim model $M_{v}$ with counterparts from the proxy model $M_{p}$. Specifically, we operate at the module- and layer-level, inspired by recent findings that malicious behaviors in LLMs are often localized to specific safety layers and safety modules (discussed in Section˜2). While not the only possible granularity, this level strikes a balance between efficiency and effectiveness, and offers greater robustness and transferability than neuron-level edits, as discussed in Section˜4.2.3 and Section˜5.3.
Our main algorithm, Guided Module-Substitution (GMS), is detailed in Section˜4.2.3. Section˜4.2.1 defines the objective optimized by GMS, and Section˜4.2.2 provides an example of constructing proxy datasets.
@@ 4.2.1 Objective
The defender’s objective is to identify a purified model $M$ within the model space $\mathbb{M}$, defined by all possible substitution strategies over the layer set $\mathcal{L}$ and module set $\mathcal{M}$, that effectively eliminates the backdoor while preserving task utility:
@@
where $\Delta_{\text{asr}}(M)$is a score that reflects the extent of backdoor removal and $\Delta_{\text{acc}}(M)$ assesses the reduction in task utility. Specifically, if given access to clean and poisoned samples, we can evaluate the attack success rate (ASR) on poisoned data and the clean accuracy (ACC) on benign inputs. The improvement in backdoor defense and degradation in utility for a given model $M$ relative to the original victim model $M_{v}$ are defined as:
@@
The hyperparameter $\alpha\in[0,1]$ controls the trade-off between backdoor removal and task utility preservation. A larger $\alpha$ favors utility, while a smaller $\alpha$ prioritizes backdoor defense.
4.2.2 Proxy datasets $\mathcal{D}_{poison}$ and $\mathcal{D}_{clean}$
To approximate Equations˜2 and 3, we extract two proxy datasets: a proxy clean set $\mathcal{D}_{\text{clean}}$ and a proxy poison set $\mathcal{D}_{\text{poison}}$, without assuming any prior knowledge of the backdoor. We only follow the standard setting to assume access to the (potentially poisoned) training set $\mathcal{D}$, or to a significantly smaller subset (e.g., 1%) of it (e.g., in user-reporting scenarios).
In literature, various methods can be used to construct proxy datasets under such settings, such as sample diagnosis Gao et al. (2019); Huang et al. (2022); Guo et al. (2023) and trigger inversion Wang et al. (2019, 2022, 2023); Xu et al. (2024). For simplicity, we adopt the following heuristics in this paper: we construct $\mathcal{D}_{\text{clean}}$ via random sampling, assuming that poisoned samples do not dominate the dataset; and we construct $\mathcal{D}_{\text{poison}}$ using a naive confidence-based heuristic, based on the observation that poisoned examples often yield higher output confidence scores on backdoored models Li et al. (2021b); Swayamdipta et al. (2020).
4.2.3 Greedy search for purified model $M_{pure}$
Using the extracted proxy datasets $\mathcal{D}_{\text{poison}}$ and $\mathcal{D}_{\text{clean}}$, we compute (potentially inaccurate) estimates of Equations˜2 and 3222An imperfect $\mathcal{D}_{\text{poison}}$ will lead to an underestimated $\Delta_{\text{asr}}(M)$, while an impure $\mathcal{D}_{\text{clean}}$ results in an overestimated $\Delta_{\text{acc}}(M)$. However, what matters is only the relative scaling between these two after weighting by $\alpha$ (Appendix C).and plug into the objective Equation˜1 to get the scoring function that guides the substitution process.
While our goal is to purify $M_{v}$ by replacing its modules across $L$ layers, an exhaustive search is computationally impractical due to its exponential complexity (e.g., $2^{6\times 12}$ for a BERT-base model (Devlin et al., 2019)). Moreover, considering individual parameters as functional units for editing often falls short in breaking backdoor connections (thereby leading prior purification works to heavily rely on clean fine-tuning to restore clean features Liu et al. (2018)), we propose a greedy algorithm that transforms the problem into a feasible search–balancing granularity and scalability by iteratively identifying and replacing the most critical module, maximizing the objective score to derive the optimal purified model $M_{pure}$.
Algorithm 1 Guided-Module-Substitution (GMS)